The hulk kicks a soccer ball with an initial velocity of 84m/s (30 degrees above horizontal)

a. What is the balls horizontal velocity?

B. What is the balls flight time?

C. What’s the balls maximum height?

D. How far forward does the ball hit the ground?

e. What’s the balls velocity when it hits the ground?
1
Expert's answer
2019-06-24T09:30:44-0400

a.


"v_x=v_0 \cos{30} =84\cos{30} =72.7\frac{m}{s}"

b.


"t=\frac{2v_0 \sin{30}}{g}=\frac{2(84) \sin{30}}{9.8}=8.6\ s"

c.


"h=\frac{(v_0 \sin{30})^2}{2g}=\frac{(84 \sin{30})^2}{2(9.8)}=90\ m"

d.


"R=\frac{(v_0)^2\sin{60}}{g}=\frac{(84)^2\sin{60}}{(9.8)}=624\ m"

e.


"v=v_0=84\frac{m}{s}"

Direction is 30 degrees below horizontal.